已知事件 AA、BB 和 CC 满足 P(A)=0.7P(A)=0.7,P(B)=0.5P(B)=0.5,P(C∣A′)=0.6P(C\mid A')=0.6 且 P(A∣C′)=0.76P(A\mid C')=0.76。

  1. (a)

    求 P(A∩B)P(A \cap B) 的最大值和最小值。

  2. (b)

    求 P(C∩A′)P(C \cap A')。

  3. (c)

    求 P(C′∩A′)P(C' \cap A')。

  4. (d)

    求 P(C)P(C)。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

7(a)

1. Find the greatest possible value of $P(A \cap B)$.

The probability of the intersection of two events cannot be greater than the probability of either individual event.
P(A∩B)≤P(A)P(A \cap B) \le P(A)
P(A∩B)≤P(B)P(A \cap B) \le P(B)
P(A∩B)≤min⁡(0.7,0.5)=0.5P(A \cap B) \le \min(0.7, 0.5) = 0.5
Therefore, the greatest possible value is 0.50.5.

2. Find the least possible value of $P(A \cap B)$.

The probability of the union of two events cannot exceed 11.
Using the addition rule, P(A∪B)=P(A)+P(B)−P(A∩B)P(A \cup B) = P(A) + P(B) - P(A \cap B)
P(A∪B)≤1P(A \cup B) \le 1
0.7+0.5−P(A∩B)≤10.7 + 0.5 - P(A \cap B) \le 1
1.2−P(A∩B)≤11.2 - P(A \cap B) \le 1
P(A∩B)≥0.2P(A \cap B) \ge 0.2
Therefore, the least possible value is 0.20.2.

答案

Greatest: 0.50.5, Least: 0.20.2

7(b)

1. Find the probability of the complement of $A$

P(A′)=1−P(A)=1−0.7=0.3P(A') = 1 - P(A) = 1 - 0.7 = 0.3

2. Use the conditional probability definition to find $P(C \cap A')$

P(C∩A′)=P(C∣A′)×P(A′)P(C \cap A') = P(C \mid A') \times P(A')
P(C∩A′)=0.6×0.3=0.18P(C \cap A') = 0.6 \times 0.3 = 0.18

答案

0.180.18

7(c)

1. Find the probabilities of the complementary events $A'$ and $C' \mid A'$.

P(A′)=1−P(A)=1−0.7=0.3P(A') = 1 - P(A) = 1 - 0.7 = 0.3
P(C′∣A′)=1−P(C∣A′)=1−0.6=0.4P(C' \mid A') = 1 - P(C \mid A') = 1 - 0.6 = 0.4

2. Calculate $P(C' \cap A')$ using the multiplication rule for conditional probability.

P(C′∩A′)=P(C′∣A′)×P(A′)P(C' \cap A') = P(C' \mid A') \times P(A')
P(C′∩A′)=0.4×0.3=0.12P(C' \cap A') = 0.4 \times 0.3 = 0.12

答案

0.120.12

7(d)

1. Find $P(A' \cap C')$ (building upon earlier parts)

P(A′)=1−P(A)=1−0.7=0.3P(A') = 1 - P(A) = 1 - 0.7 = 0.3

P(A′∩C′)=P(C′∣A′)P(A′)=(1−0.6)×0.3=0.12P(A' \cap C') = P(C' \mid A') P(A') = (1 - 0.6) \times 0.3 = 0.12

2. Set up an equation for $P(C')$ using the definition of conditional probability

Since C′C' is partitioned into mutually exclusive events (A∩C′)(A \cap C') and (A′∩C′)(A' \cap C'), we have:
P(A∩C′)=P(C′)−P(A′∩C′)=P(C′)−0.12P(A \cap C') = P(C') - P(A' \cap C') = P(C') - 0.12

Using P(A∣C′)=0.76P(A \mid C') = 0.76:
P(A∩C′)P(C′)=0.76\frac{P(A \cap C')}{P(C')} = 0.76
P(C′)−0.12P(C′)=0.76\frac{P(C') - 0.12}{P(C')} = 0.76

3. Solve for $P(C')$ and hence find $P(C)$

1−0.12P(C′)=0.761 - \frac{0.12}{P(C')} = 0.76

0.12P(C′)=0.24\frac{0.12}{P(C')} = 0.24

P(C′)=0.120.24=0.5P(C') = \frac{0.12}{0.24} = 0.5

Therefore, P(C)=1−P(C′)=1−0.5=0.5P(C) = 1 - P(C') = 1 - 0.5 = 0.5

答案

0.50.5

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  4. 4已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  5. 5对于事件 AA 和 BB,已知 P(A)=0.55\mathrm{P}(A) = 0.55,P(B)=0.65\mathrm{P}(B) = 0.65 且 P(A∪B)=0.93\mathrm{P}(A \cup B) = 0.93。 求 P(A∩B)\mathrm{P}(A \cap B)…